OpenStudy (mgriffin16):

Write the equation of the following statement: 4 times the sum of a number and 3 is 9 less than twice the number. A. 4(b + 3) = 9 - 2b B. 4b + 3 = 2b - 9 C. 4(b + 3) = 2b - 9 D. 4(b - 3) = 9 - 2b ????

OpenStudy (okdutchman7):

4(N+3) +9 = 2N

OpenStudy (okdutchman7):

Can you figure it out if I give you that?

OpenStudy (okdutchman7):

Just rearrange the equation I gave you and use b instead of n

OpenStudy (0487308):

First you have to isolate N. Do you know how to do that, or do you need me to help you with that?

OpenStudy (0487308):

Sorry, I mean N. I was looking at okdutchman7's reply XD

OpenStudy (okdutchman7):

4(n+3) = 2n - 9

OpenStudy (mgriffin16):

its just asking me to put that sentence into a equation, i do not have to solve anything. It will either be A,B,C, or D

OpenStudy (okdutchman7):

Change what I wrote from n to b and that is your answer

OpenStudy (okdutchman7):

I can't give you the answer directly.

OpenStudy (mgriffin16):

oh thank you

OpenStudy (0487308):

Okay. First, you will have to subtract 9 from each side. This will give you 4(b+3) = 2b-9.

OpenStudy (0487308):

The answer is B.

OpenStudy (mgriffin16):

Thank you very much

OpenStudy (0487308):

Okay, no problem. Do you understand now?